OpenStudy (anonymous):

simplify: 42a-18b -------- 6

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge \frac{42a-18b}{6}=\frac{1}{6}*(42a-18b)$$ $$\Huge \frac{x}{y}=\frac{1}{y}*x$$

OpenStudy (anonymous):

yup now take 6 common out from numerator and forms the last answer .

OpenStudy (anonymous):

what is its simplest form?

OpenStudy (anonymous):

and thanks

OpenStudy (skullpatrol):

Use the distributive property to simply. $$\Huge a(b+c)=ab+ac$$

OpenStudy (anonymous):

okay then then with the number the answer would be 7a+18b?

OpenStudy (skullpatrol):